### v2

# thanks: The first two authors contributed equally

Abstract
The Inspiral Merger Ringdown Consistency Test (IMRCT) is one among a battery of tests of general relativity (GR) employed by the LIGO-Virgo-KAGRA (LVK) collaboration. It is used to search for deviations from GR in detected gravitational waves (GWs) from compact binary coalescences (CBCs) in a model-agnostic way. The test compares source parameter estimates extracted independently from the inspiral and post-inspiral portions of the CBC signals and, therefore, crucially relies on the accurate modeling of the waveform. Current implementations of the IMRCT routinely use quasicircular waveforms, under the assumption that the residual eccentricity of the binary when the emitted GWs enter the frequency band of the LVK detector network will be negligible. In this work, we perform a detailed study to investigate the typical magnitudes of this residual eccentricity that could potentially lead to spurious violations of the IMRCT. To that end, we conduct injection campaigns for a range of eccentricities and recover with both quasicircular and eccentric waveforms. We find that an eccentric GW signal from a GW150914-like system with eccentricity $e_{\text{gw}}\gtrsim$ 0.04 at an orbit averaged frequency $\langle f_{\text{ref}}\rangle=$ 25 Hz breaks the IMRCT if recovered with quasicircular waveforms at $\gtrsim 68\%$ confidence. The violation becomes more severe ($\gtrsim 90\%$ confidence) for $e_{\text{gw}}=$ 0.055 at $\langle f_{\text{ref}}\rangle=$ 25 Hz. On the other hand, when eccentric waveforms are used, the IMRCT remains intact for all eccentricities considered. We also briefly investigate the effect of the magnitude and orientation (aligned/antialigned) of the component spins of the binary on the extent of the spurious violations of the IMRCT. Our work, therefore, demonstrates the need for accurate eccentric waveform models in the context of tests of GR.
I Introduction
The LIGO-Virgo Aasi et al. (2015); Acernese et al. (2015) network of ground-based interferometric gravitational-wave (GW) detectors has detected $\sim 90$ compact binary coalescence (CBC) events in its first three observing runs (O1, O2, O3) Abbott et al. (2023). The majority of these are binary black hole (BBH) mergers, although binary neutron star (BNS) Abbott et al. (2017, 2020a) and neutron star-black hole (NSBH) coalescences Abbott et al. (2021a) have also been observed.
The detected GW signals, especially those pertaining to BBH mergers, come from the last few orbits of the inspiral, as well as the merger and ringdown phases (see, e.g., Sathyaprakash and Schutz (2009)). These GWs are well-suited to probe general relativity (GR) in the strong-field regime, unlike most tests of GR that use electromagnetic waves, which typically probe the weak field regime [^1] A notable exception to this are the results of the Event Horizon Telescope (EHT) Akiyama et al. (2019), which imaged the shadow of the supermassive black hole at the center of the M31 galaxy and enabled electromagnetic-based strong-field tests of GR Psaltis (2019)..
The LIGO-Virgo-KAGRA (LVK) collaboration  Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) has conducted a suite of tests of GR across O1, O2, and O3 Abbott et al. (2021b). This includes222but is not limited to
  • A model-agnostic residuals test, which subtracts out the best-matched GR-modeled CBC waveform from the data containing the GW signal, and checks if the statistical properties of the residual are consistent with noise Abbott et al. (2016, 2019a, 2020b)
  • A test that probes the inspiral evolution of the CBC by searching for deviations in Post-Newtonian (PN) parameters of the GW signal Blanchet and Sathyaprakash (1995, 1994); Arun et al. (2006a, b); Yunes and Pretorius (2009); Mishra et al. (2010); Li et al. (2012a, b)
  • An inspiral-merger-ringdown consistency test (IMRCT) that compares and assesses the consistency between the low and high-frequency portions of the CBC signals Ghosh et al. (2016, 2018)
  • Propagation tests of GWs that compare their speed with respect to the speed of light Abbott et al. (2019b) as well as any modulations in the waveform due to finite-graviton-mass-driven velocity dispersion Will (1998).
In this work, we restrict our attention to the IMRCT. The IMRCT compares the two-dimensional joint posterior on the final mass and spin of the merged binary estimated independently from the inspiral (low-frequency) and post-inspiral (high-frequency) portions of the waveform Ghosh et al. (2016, 2018). A sufficiently large deviation from GR should result in the difference distribution between final mass and final spin evaluated from these 2D posteriors to deviate from the origin (zero). However, it is conceivable that waveforms that do not accurately capture the physics of the CBC could lead to biased posteriors. We propose to study the resulting systematics in the IMRCT. In particular, we investigate how the neglect of eccentricity in the waveform models could cause spurious violations of this test.
It is well-known that, as GWs carry away energy and angular momentum of a CBC, it reduces the binary orbit’s eccentricity Peters and Mathews (1963); Peters (1964). The prevalent expectation is that by the time GWs from the CBC enter the frequency band of the LVK network, any initial eccentricity that the binary may have had at the time of formation would have been reduced to a negligibly small value. Indeed, current constraints on the rate of eccentric mergers, from searches for such signals, suggest that observations of these in O4 are likely to be relatively rare Abac et al. (2023); Abbott et al. (2019c); Nitz et al. (2019). Thus, tests of GR conducted so far by the LVK routinely use quasicircular waveforms, including the IMRCT.
On the other hand, certain formation channels, such as those pertaining to dense stellar environments, could produce binaries whose eccentricities are non-negligible when they enter the LVK frequency band Mapelli (2020). Dynamical encounters in such environments could harden the binary so that GW emission cannot completely carry away the eccentricity by the time the GWs enter the LVK’s frequency band Samsing (2018); Zevin et al. (2019); Romero-Shaw et al. (2022). Moreover, in active-galactic-nuclei disks, eccentricity can be amplified by binary-single encounters Samsing et al. (2022); Tagawa et al. (2021), while the presence of a third object in the vicinity of the binary could similarly boost the eccentricity via the Kozai-Lidov mechanism Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Randall and Xianyu (2018); Yu et al. (2020); Bartos et al. (2023). Such eccentricity-enhancing processes could also leave a residual in-band eccentricity detectable by the LVK network. Neglecting eccentricity of such signals in (PN) parametrized tests has already been shown to lead to spurious violations of GR both at the individual event level Saini et al. (2022) and population level inference Saini et al. (2023). We here focus only on the effect of neglect of eccentricity on the IMRCT.
Performing the IMRCT using quasicircular waveform models, could potentially cause systematics in the IMRCT. Intuitively, the eccentricity at the time of band-entry of the GW signal will be larger than the eccentricity in the post-inspiral phase, where it is usually vanishingly small. Thus, one would expect source parameters recovered with quasicircular waveform models to be biased when using the inspiral portion of the waveform Favata (2014); Favata et al. (2022); O’Shea and Kumar (2023), while those recovered using the post-inspiral phase to be relatively less biased, if at all. The resulting 2D difference-distribution constructed from the (inspiral/post-inspiral) final mass and spin posteriors should be deviated from zero.
To assess the impact of neglecting eccentricity in waveform models on the IMRCT, we consider a set of synthetic eccentric GW signals. We inject these in the most probable realization of (zero-mean) Gaussian noise, the so-called “zero-noise” realization [^3] Here zero-noise realization refers to the most-probable realization of Gaussian noise with zero mean and the parameter uncertainties obtained using this noise-realization provide lower bound to the estimates obtained using the non-zero real noise-realization., assuming an O4-like noise power-spectral density (PSD) Abbott et al. (2018). We then perform the IMRCT on each of these signals, using quasicircular templates. We find that the IMRCT is indeed violated with quasicircular templates, as we increase the injected eccentricity, for a set of masses and spins where the IMRCT is known not to suffer from other systematics. This violation progressively increases, from moderate to severe, with increasing eccentricity. To ascertain that no intrinsic waveform systematics are contributing to this bias, we repeat the IMRCT using eccentric waveform model for recovery. We indeed find that the IMRCT is no longer violated.
Past work, using an approximate Fisher Matrix Analysis (FMA) Cutler and Flanagan (1994) and the Cutler-Vallisneri framework Cutler and Vallisneri (2007), pointed out that the IMRCT could be violated due to neglect of eccentricity Bhat et al. (2023). Another work also found spurious violations of the IMRCT using available numerical-relativity waveforms as injections, while performing the IMRCT with quasicircular phenomenological waveform models Narayan et al. (2023). Our work complements and expands on both of these studies in the following way. We perform Bayesian parameter inference runs using large-scale nested sampler and, therefore, have more realistic estimate of the bias compared to those in provided in Bhat et al. (2023). Moreover, unlike in Bhat et al. (2023), we do not neglect the eccentricity in the post-inspiral part of the signal which may lead to over-estimation or under-estimation of the bias. Also, because we employ an eccentric waveform model TEOBResumS-Dalí Nagar et al. (2018, 2021a), we have two advantages over the results in  Narayan et al. (2023). We use the same waveform model to create the injection and then to recover it. Therefore, we rule out any bias originating due to the waveform systematics which may be present if the recovery template is not as accurate as the injected NR waveforms. Also, we have the freedom to create injections at arbitrary values of eccentricities allowing us to do an injection study on a finer grid of eccentricity values. This, in turn, allows us to identify more accurately at what eccentricity the IMRCT bias becomes severe. Similarly, we examine the sensitivity of the bias for a given eccentricity on other parameters like the magnitude of the component spins and their orientation (aligned/antialigned) with respect to the orbital angular momentum of the binary. Moreover, we ascertain that the IMRCT is indeed recovered as expected, when eccentricity is included, for all eccentricities considered. We make contact with both these works to compare the eccentricity at which the IMRCT is violated, using a generalised definition of eccentricity that is independent of waveform models Shaikh et al. (2023).
The rest of this paper is organized as follows. Section II briefly describes the IMRCT, and argues why one should expect spurious violations due to neglect of eccentricity. Section III shares the results of the injection campaign described above. Section IV summarizes the work and discusses other potential sources of spurious IMRCT violations.
II Methods
II.1 Gravitational-wave parameter estimation
In GR, GWs have two polarizations, $h_{+}$ and $h_{\times}$. The phase evolution of these polarizations depends on the intrinsic parameters of the binary, viz. the masses $m_{1},m_{2}$ of the individual components  [^4] Strictly, the masses that determine the shape of the GW signal measured at the detector are the so-called “detector-frame” masses, which are simply the masses multiplied by $(1+z)$, where $z$ is the cosmological redshift., and the spin vectors of these components $\mathbf{S}_{1},\mathbf{S}_{2}$. The amplitude of the polarizations depend on the inclination of the binary with respect to the line of sight, $\iota$, luminosity distance $d_{L}$, the phase $\phi_{c}$ and time $t_{c}$ at coalescence. In addition to the above mentioned intrinsic parameters, we need two more parameters – eccentricity $e$ and mean anomaly $l$  [^5] The angle variable that describes the position of the compact objects on the orbit can be also described by other equivalent variables like eccentric anomaly or true anomaly. Clarke et al. (2022) – at a given reference frequency to accurately describe GWs from a binary in eccentric orbit. Also, the GW waveform shape is more sensitive to the ‘chirpmass’ $\mathcal{M}_{c}=m_{1}^{3/5}m_{2}^{3/5}/(m_{1}+m_{2})^{1/5}$ compared to individual masses and often it is more practical to use $\mathcal{M}_{c}$ and the mass ratio $q=m_{1}/m_{2}$ instead of $m_{1}$ and $m_{2}$ for parameter estimation (PE) Cutler and Flanagan (1994).
The response of the GW detector to each of the GW polarizations is given by the antenna pattern functions, $F_{+},F_{\times}$. These depend on the location of the source in the sky, viz. the right ascension $\alpha$ and declination $\delta$, as well as the polarization angle $\psi$. The GW strain measured by the detector is a linear combination of the two polarizations, weighted by the antenna pattern functions:
$h(t;\vec{\theta},\vec{\lambda})=F_{+}(\vec{\lambda}_{F})h_{+}(t;\vec{\theta},\vec{\lambda}_{h})+F_{\times}(\vec{\lambda}_{F})h_{\times}(t;\vec{\theta},\vec{\lambda}_{h})$ (1)
with $t$ as the detector-frame time, $\vec{\theta}=\{{\mathcal{M}_{c},q,\mathbf{S}_{1},\mathbf{S}_{2},e,l\}}$, $\vec{\lambda}_{F}=\{{\alpha,\delta,\psi\}}$, $\vec{\lambda}_{h}=\{{d_{L},\iota,t_{c},\phi_{c}\}}$, and $\vec{\lambda}=\vec{\lambda}_{F}\cup\vec{\lambda}_{h}$.
Given two waveforms $h_{1,2}(t;\vec{\theta},\vec{\lambda})$, their noise-weighted inner product is defined as:
[…]
where $f$ is the frequency, and $\tilde{h}_{1,2}$ are Fourier transforms of the corresponding time domain waveforms, and $\star$ represents complex conjugation and $\Re$ denotes the real part. $S_{n}(f)$ is the noise power spectral density (PSD) of the detector. The lower and the upper cutoff frequencies, $f_{\text{min}}$ and $f_{\text{max}}$ are determined by the source properties and the sensitivity of the detector.
Let us now consider a time-domain stretch of detector strain data:
$s(t)=n(t)+h(t;\vec{\theta},\vec{\lambda})$ (3)
known to consist of noise $n(t)$, and GW signal $h(t;\vec{\theta},\vec{\lambda})$ whose parameters $\vec{\theta},\vec{\lambda}$ are to be estimated. This can be achieved by sampling the seventeen-dimensional GW posterior $p(\vec{\theta},\vec{\lambda}|s)\propto p(\vec{\theta},\vec{\lambda})p(s|\vec{\theta},\vec{\lambda})$, where $p(\vec{\theta},\vec{\lambda})$ is the prior distribution on the parameters, and the GW likelihood is given by Cutler and Flanagan (1994):
$p(s|\vec{\theta},\vec{\lambda})\propto\exp\left[-\frac{\langle s-h(\vec{\theta},\vec{\lambda}),s-h(\vec{\theta},\vec{\lambda})\rangle}{2}\right]$ (4)
The sampling of this high-dimensional posterior is typically achieved using large-scale Markov-chain Monte-Carlo methods, or nested-sampling methods.
II.2 The inspiral-merger-ringdown consistency test
The IMRCT can be thought of as a coarse-grained test of the frequency evolution of a BBH waveform. It compares the low-frequency and high-frequency portions of the waveform by comparing the final mass and spin of the merged binary estimated from the inspiral and merger-ringdown regimes. Let $M_{f}^{\text{I}}$ and $\chi_{f}^{\text{I}}\equiv|S_{f}^{\text{I}}|/(M_{f}^{\text{I}})^{2}$ be the estimates of the detector-frame final mass and dimensionless spin of the merged binary, respectively, from the inspiral and $M_{f}^{\text{MR}}$, $\chi_{f}^{\text{MR}}$ be the same from the post-inspiral part. The IMRCT constructs a joint distribution on the following difference parameters Ghosh et al. (2016, 2018):
$\frac{\Delta M_{f}}{\bar{M}_{f}}\equiv 2\frac{M_{f}^{\text{I}}-M_{f}^{\text{MR}}}{M_{f}^{\text{I}}+M_{f}^{\text{MR}}},\leavevmode\nobreak\ \frac{\Delta\chi_{f}}{\bar{\chi}_{f}}\equiv 2\frac{\chi_{f}^{\text{I}}-\chi_{f}^{\text{MR}}}{\chi_{f}^{\text{I}}+\chi_{f}^{\text{MR}}}$ (5)
The joint-distribution on these parameters should be consistent with zero if the waveform model used to extract the signal in the data captures all the physics associated with the production of the signal. On the other hand, a sufficiently large deviation from GR is expected to shift this distribution to an extent that it does not contain zero with high confidence ($\gtrsim 68\%$ or $\gtrsim 90\%$). In this work, we demonstrate that even if both the waveform model and the signal in the data are consistent with GR, an eccentric signal recovered with a quasicircular model will also lead to a (spurious) violation of the IMRCT.
The estimate of the final mass and final spin of the merged binary determined with the intrinsic parameters $\vec{\theta}$ extracted from the inspiral portion of the waveform is achieved via numerical-relativity-based fits Hofmann et al. (2016); Healy and Lousto (2017); Jiménez-Forteza et al. (2017), as done in Ghosh et al. (2016, 2018). On the other hand, the shape of the ringdown contained in the post-inspiral portion of the waveform depends on the final mass and spin, which can therefore be extracted directly.
In this work, we restrict ourselves to non-precessing waveforms, i.e, waveforms produced by binaries whose spin-vectors are aligned or antialigned with the orbital angular momentum vector $\mathbf{L}$ ($z$-axis). Moreover, we only consider binaries with mass-ratios $q$ that are close to 1 ($q=m_{1}/m_{2}\sim 1$). We therefore only need to use the leading order $(2,\pm 2)$ multipole, since higher multipoles for the types of systems we consider will be suppressed.
The choice of the transition frequency that separates the inspiral and post-inspiral portions of the waveform is not unique. However, all implementations of the IMRCT use some approximation of the frequency corresponding to the innermost stable circular orbit (ISCO) of the binary. The LVK estimates the ISCO using the medians of intrinsic parameters $\theta$ extracted from the full inspiral-merger-ringdown (IMR) waveform. Narayan et al Narayan et al. (2023) use the medians of the final mass and spin of the merged binary estimated from the full IMR waveform, similar to what was originally done in Ghosh et al. (2016, 2018). However, as pointed out in Ghosh et al. (2016, 2018); Narayan et al. (2023), the IMRCT is largely insensitive to moderate variations in the choice of the transition frequency. In this work, we use the exact values of $\vec{\theta}$ (i.e, the injected values), as well as the numerical relativity fits Hofmann et al. (2016); Healy and Lousto (2017); Jiménez-Forteza et al. (2017), to determine a point-estimate of the final mass and spin, from which the ISCO frequency is evaluated.
II.3 A standardized definition of eccentricity
In GR, there is no universally agreed-upon definition of eccentricity. Consequently, different waveform models adopt custom internal definitions of eccentricity, leading to model-dependent inferences about eccentricity. To mitigate ambiguity in the inferred value of eccentricity from gravitational wave data analysis, there is a growing effort to standardize the definition of eccentricity. This standardization involves measuring eccentricity from a gauge-independent quantity, such as the gravitational waveform modes Mora and Will (2002); Ramos-Buades et al. (2020); Islam et al. (2021); Nagar et al. (2021a); Ramos-Buades et al. (2022a); Bonino et al. (2023); Ramos-Buades et al. (2022b); Shaikh et al. (2023).
The waveform modes are obtained from the polarizations by decomposing the complex combination $h_{+}-ih_{\!\times\!}$ on a sphere into a sum of spin-weighted spherical harmonic modes $\mathpzc{h}_{\ell m}$, so that the waveform along any direction $(\iota,\varphi_{0})$ in the binary’s source frame is given by
$\mathpzc{h}(t,\iota,\varphi_{0})=\sum_{\ell=2}^{\ell=\infty}\sum_{m=-\ell}^{m=\ell}\mathpzc{h}_{\ell m}(t)\leavevmode\nobreak\ {}_{-2}Y_{\ell m}(\iota,\,\varphi_{0}),$ (6)
where $\iota$ and $\varphi_{0}$ are the polar and azimuthal angles on the sky in the source frame, and ${}_{-2}Y_{\ell m}$ are the spin $=-2$ weighted spherical harmonics.
In this work, we employ the Python implementation gw_eccentricity of the standardized definition of eccentricity from Shaikh et al. (2023), which defines the eccentricity $e_{\text{gw}}$ using the frequency of the (2, 2) mode $\mathpzc{h}_{22}$. Initially, it computes $e_{\omega_{22}}$ as described in Ramos-Buades et al. (2020); Islam et al. (2021); Bonino et al. (2023); Shaikh et al. (2023).
$e_{\omega_{22}}(t)=\frac{\sqrt{\omega^{\text{p}}_{22}(t)}-\sqrt{\omega^{\text{a}}_{22}(t)}}{\sqrt{\omega^{\text{p}}_{22}(t)}+\sqrt{\omega^{\text{a}}_{22}(t)}},$ (7)
Here, $\omega^{\text{p}}_{22}(t)$ and $\omega^{\text{a}}_{22}(t)$ denote the interpolants through the frequency of the (2, 2) mode at the pericenter (point of closest approach) and the apocenter (point of farthest approach), respectively. The frequency $\omega_{22}$ of the (2, 2) mode is derived from $\mathpzc{h}_{22}$ as
[…]
$\omega_{22}(t)=\frac{\mathrm{d}\phi_{22}(t)}{\mathrm{d}t}\,.$ (9)
After computing $e_{\omega_{22}}$ from $\mathpzc{h}_{22}$, $e_{\text{gw}}$ is determined through the following transformation to ensure that $e_{\text{gw}}$ correctly exhibits the Newtonian limit Ramos-Buades et al. (2022b).
$e_{\text{gw}}=\cos(\Psi/3)-\sqrt{3}\,\sin(\Psi/3),$ (10)
where
$\Psi=\arctan\left(\frac{1-e_{\omega_{22}}^{2}}{2\,e_{\omega_{22}}}\right).$ (11)
Since eccentricity decays via gravitational wave radiation, we must specify a reference frequency where the value of $e_{\text{gw}}$ is measured. Because eccentricity induces modulations in the instantaneous frequency $\omega_{22}$, we obtain a monotonic orbit-averaged $\langle\omega_{22}\rangle$ by taking the orbital average of the instantaneous $\omega_{22}$ (see Section II.E.2 of Shaikh et al. (2023) for more details) Between any two consecutive extrema (local maxima (pericenters, $t^{\text{p}}$) or minima (apocenters, $t^{\text{a}}$) in the instantaneous $\omega_{22}$) $t^{\text{X}}_{i}$ and $t^{\text{X}}_{i+1}$ one can define
$\langle\omega_{22}\rangle_{i}^{\text{X}}$ $=\frac{1}{t^{\text{X}}_{i+1}-t^{\text{X}}_{i}}\,\int_{t^{\text{X}}_{i}}^{t^{\text{X}}_{i+1}}\omega_{22}(t)\,\mathrm{d}t$
$=\frac{\phi_{22}(t^{\text{X}}_{i+1})-\phi_{22}(t^{\text{X}}_{i})}{t^{\text{X}}_{i+1}-t^{\text{X}}_{i}},$ (12)
and associate $\langle\omega_{22}\rangle_{i}^{\text{X}}$ with the midpoint between $t^{\text{X}}_{i}$ and $t^{\text{X}}_{i+1}$:
$\langle t\rangle^{X}_{i}=\frac{1}{2}\left(t^{\text{X}}_{i}+t^{\text{X}}_{i+1}\right).$ (13)
Applying this procedure to all consecutive pairs of pericenter times, one can obtain the set $\big{\{}\,\left(\langle t\rangle^{\text{p}}_{i},\langle\omega_{22}\rangle_{i}^{\text{p}}\right)\,\big{\}}$ and similarly, the set $\left\{\,\big{(}\langle t\rangle^{\text{a}}_{i},\langle\omega_{22}\rangle_{i}^{\text{a}}\big{)}\,\right\}$ for apocenter times. Finally, taking the union of these two datasets, one can build an interpolant in time to obtain $\langle\omega_{22}\rangle(t)$ and use $\langle f_{\text{ref}}\rangle(t_{\text{ref}})\equiv\langle\omega_{22}\rangle(t_{\text{ref}})/(2\pi)$ for defining the reference frequency $\langle f_{\text{ref}}\rangle$ corresponding to a reference time $t_{\text{ref}}$.
Presenting our results with a standardized definition facilitates one-to-one comparisons with similar studies in the past and enables future studies to do the same with the results in this work.
II.4 Parameter Estimation setup
We obtain our results by performing Bayesian parameter inference runs using state-of-the-art GW parameter inference framework built in Bilby Ashton et al. (2019) with a dynamical nested sampler dynesty Speagle (2020). For the injection and recovery, we use the publicly available eccentric waveform model TEOBResumS-Dalí Nagar et al. (2018, 2021b); Chiaramello and Nagar (2020); Albanesi et al. (2021); Nagar et al. (2021a); Nagar and Rettegno (2021); Nagar et al. (2024)  [^6] We use the commit 0f19532 of the eccentric branch of TEOBResumS available at https://bitbucket.org/eob_ihes/teobresums/. In the following, we will denote the eccentricity that the waveform model takes as an input by $e_{\text{eob}}$ and the frequency where $e_{\text{eob}}$ is defined by $f_{\text{ref}}$. Additionally, we provide $\langle f_{\text{ref}}\rangle,e_{\text{gw}}$} for each injection to facilitate better comparisons across waveform models and systematics studies.  [^7] Note that we sample over the model eccentricity $e_{\text{eob}}$ and not $e_{\text{gw}}$. For a given model eccentricity $e_{\text{eob}}$ at a model reference frequency $f_{\text{ref}}$, the standardized eccentricity $e_{\text{gw}}$ at $\langle f_{\text{ref}}\rangle$ is obtained by first generating the eccentric waveform modes for the given pair $\{f_{\text{ref}},e_{\text{eob}}\}$ and then following the steps discussed in Sec. II.3.
Ideally, to accurately describe an eccentric orbit, the waveform model should also include mean anomaly as a free parameter Islam et al. (2021); Clarke et al. (2022); Shaikh et al. (2023). In this work, we use the version of TEOBResumS-Dalí that was used in Ref. Bonino et al. (2023) by the waveform developers to showcase the accuracy and robustness of their model. This version of TEOBResumS-Dalí does not allow mean anomaly as a free parameter. Although the most recent version of TEOBResumS-Dalí, released after the completion of this work, allows mean anomaly as a free parameter, this version is yet to be tested (at the time of writing) in a parameter inference study by sampling over mean anomaly and is currently under review. Eccentric waveform models Ecc1dSur Islam et al. (2021) and SEOBNRv4EHM Ramos-Buades et al. (2022a, 2023) include mean anomaly as a free parameter but these are not publicly available. Publicly available eccentric waveform models SEOBNRE Cao and Han (2017); Liu et al. (2020) and EccentricTD Tanay et al. (2016) do not include mean anomaly as a free parameter but instead start at the pericenter or apocenter. [^8] In some classes of GW sources, eccentricity and mean anomaly may not be sufficient to describe the orbit. For example, in Ref. Carullo et al. (2024), the Arnowitt-Deser-Misner energy and angular momentum are used to describe a generic planar orbit. See also Ref. Gamba et al. (2023) which use the same parameterization to study the possibility of GW190521 resulting from a dynamical capture of two nonspinning black holes.
Because the version of TEOBResumS-Dalí used in the work does not allow sampling over mean anomaly, Ref. Bonino et al. (2023) sampled over the reference frequency, i.e., the PE was performed by sampling over eccentricity and reference frequency with fixed mean anomaly. The motivation behind such an approach is that by varying the reference frequency with fixed mean anomaly one can expect to explore the same parameter space as in the case of varying mean anomaly at a fixed reference frequency. For example, in a recent work in Ref. Bonino et al. (2024), it was shown that one can map an eccentric Numerical Relativity (NR) waveform to an eccentric effective one body waveform using just the initial eccentricity and reference frequency at the first apocenter (that is, at a fixed mean anomaly). Although sampling over reference frequency may reduce bias in PE compared to a fixed reference frequency when mean anomaly is fixed, no significant bias was found in the recovered parameters when sampling exclusively over eccentricity with both mean anomaly and reference frequency fixed (see Fig. 3. in Ref. Bonino et al. (2023)).
Contrary to Ref. Bonino et al. (2023, 2024) which suggest that using (eccentricity, mean anomaly) at fixed reference frequency and using (eccentricity, reference frequency) at fixed mean anomaly is equivalent, it has also been argued that sampling over eccentricity and reference frequency with fixed mean anomaly may not be equivalent to sampling over eccentricity and mean anomaly with a fixed reference frequency (see the introduction in Ref. Gupte et al. (2024)) since these two methods may not explore the same parameter space. In addition, varying reference frequency changes the reference point where binary properties are measured.
Given the fact that no significant bias was reported by the waveform developers in Ref. Bonino et al. (2023) when sampling only eccentricity with fixed reference frequency and that this version does not allow varying mean anomaly, we restrict our PE to sample only over eccentricity. However, future works that will use the recent version of TEOBResumS-Dalí once it is reviewed, should sample over mean anomaly to reduce any potential bias due to neglecting the mean anomaly.
To obtain the 2D posteriors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ for the IMRCT, we use the summarytgr command line tool from PESummary Hoy and Raymond (2021). It provides the posteriors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ in a post-processing step making use of mass and spin posterior samples along with the NR-fits for the final mass and the final spin parameters. The uninformative priors used for the mass and spin parameters force non-trivial undesirable priors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ obtained in the post-processing. To revoke this effect and to impose agnostic priors on these inference parameters used to probe GR-violation, we reweight [^9] Reweighting of posterior refers to removing the effect of any non-trivial undesirable prior on the posterior (see Chapter 5 in Ref. Gelman (2013) for details). This is done when there is no prior information about the parameter in consideration or the prior information about the parameter is unphysical and one wants to be as objective as possible. the 2D posterior to a flat prior on the deviation parameters, as is routinely done by the LVK Abbott et al. (2021c, b).
Given the posterior samples obtained in the PE from the inspiral and post-inspiral part, summarytgr uses an NR informed fit to obtain the final mass and spin posteriors. This fit is based on quasicircular NR simulations only and does not take into account any eccentric corrections which has been shown to be negligible at the quasi-Newtonian level with small eccentricity approximation Bhat et al. (2023). Currently, we lack a sufficient number of NR-simulations of merging eccentric binaries required for obtaining the eccentricity corrected NR-fit for final mass and final spin. We measure $\langle f_{\text{ref}}\rangle,e_{\text{gw}}$} for a given {$f_{\text{ref}},e_{\text{eob}}$} using gw_eccentricity Shaikh et al. (2023).
III Results
The main result of this work can be divided into two parts. The first part, Sec. III.1, discusses the effect of neglecting eccentricity in the waveform template on the IMRCT. The second part, Sec. III.2, demonstrates the robustness of the IMRCT performed on GW data containing an eccentric signal using a waveform template that includes physics required to describe GWs coming from binaries on an eccentric orbit. We list the injection values and priors in 1 and summarize the results of the study of bias in IMRCT due to neglecting eccentricity in 2.
III.1 Effects of neglecting eccentricity in templates on IMRCT
Figure 1: Bias in IMRCT due to neglecting eccentricity in recovery template. The top and the bottom row represent the IMRCT result for the aligned spin configurations with dimensionless component spins $(\chi_{1},\chi_{2})=(0.4,0.3)$ and $(\chi_{1},\chi_{2})=(0.2,0.1)$, respectively. The contours represent the joint probability distribution of the final mass deviation $\Delta M_{f}/\bar{M}_{f}$ and final spin deviation $\Delta\chi_{f}/\bar{\chi}_{f}$ at 90% and 68% confidence on the left and the right panels, respectively. For $(\chi_{1},\chi_{2})=(0.4,0.3)$, GR value is excluded at 68% confidence at $e_{\text{eob}}=0.09$. For $(\chi_{1},\chi_{2})=(0.2,0.1)$, GR value is excluded at 68% and 90% confidence for $e_{\text{eob}}=0.05$ and $0.07$, respectively.
III.1.1 Aligned spin case
First, we focus on studying the cases where the spins of the binary components are aligned with the orbital angular momentum of the binary. It is convenient to define the tilts $\theta^{\mathbf{S}\mathbf{L}}_{i}$ as the angle between of the component spin $\mathbf{S}_{i}$ and the orbital angular momentum $\mathbf{L}$
$\theta^{\mathbf{S}\mathbf{L}}_{i}=\cos^{-1}\left(\frac{\mathbf{L}.\mathbf{S}_{i}}{|\mathbf{L}||\mathbf{S}_{i}|}\right)\qquad i=\{1,2\}$ (14)
Since TEOBResumS-Dalí can describe only non-precessing systems, $\theta^{\mathbf{S}\mathbf{L}}_{i}$ can take only two values 00 (aligned) or $\pi$ (antialigned). In this section, we consider binary systems with $\theta^{\mathbf{S}\mathbf{L}}_{1,2}=0$. Since the waveform model only allows tilts of either 0 or $\pi$, and does not allow sampling over arbitrary spin orientations, we choose the prior on the tilts to be a Dirac delta function peaked at the injection value.
We choose the component masses to be similar to that of the gravitational wave event GW150914. For this study we choose a total mass of system that lies within the ‘golden mass’ range $M\in[50-100]M_{\odot}$ which ensures high enough SNRs both in the inspiral and post-inspiral parts of the signal Ghosh et al. (2016). The injection parameters and the priors are noted in Table 1. The injection eccentricities $e_{\text{eob}}\in$0.05, 0.06, 0.07, 0.08, 0.09, 0.1} are defined at a reference frequency of $f_{\text{ref}}=$ 20 Hz. These injected waveforms have $e_{\text{gw}}$ values {0.040, 0.048, 0.055, 0.063, 0.071, 0.079}  [^10] We used the ResidualAmplitude method within gw_eccentricity to measure these eccentricities. measured at an orbit averaged frequency of $\langle f_{\text{ref}}\rangle=$ 25 Hz. To understand the effect of spin magnitudes on IMRCT systematics, we consider two different choices of dimensionless component spins, one low $(\chi_{1},\chi_{2})=(0.2,0.1)$ and one moderate $(\chi_{1},\chi_{2})=(0.4,0.3)$, for the set of injected eccentricities. All the injections considered in this study have Hanford-Livingston-Virgo network (at O4 design sensitivity) signal to noise ratio (SNR) in the range $27-52$.
While recovering the injection, we use a Dirac delta prior on eccentricity peaked at $e_{\text{eob}}=10^{-4}$ making the parameter estimation procedure equivalent to one using a quasicircular template.  [^11] For $e_{\text{eob}}\leq 10^{-4}$, the corresponding $e_{\text{gw}}$ is $10^{-6}$ implying the waveforms to be essentially quasicircular. See fig. 10 in  Shaikh et al. (2023)
Parameter Injected value Prior
$\mathcal{M}_{c}(M_{\odot})$ 26.36 $\mathcal{U}(12,45)$
$q$ 0.8 $\mathcal{U}(0.25,1)$
$\chi_{1}$0.2, 0.4} $\mathcal{U}(0,0.99)$
$\chi_{2}$0.1, 0.3} $\mathcal{U}(0,0.99)$
$\theta^{\mathbf{S}\mathbf{L}}_{1}$ (rad) 0 0
$\theta^{\mathbf{S}\mathbf{L}}_{2}$ (rad) 0 0
$d_{L}$(MPc) 410 $\mathcal{U}(50,2000)$
$\iota$ $0.78$ $\mathcal{\sin}(0,\pi)$
$\alpha$(rad) 4.89 $\mathcal{U}(0,2\pi)$
$\delta$0.218 $\mathcal{\cos}(-\pi/2,\pi/2)$
$\psi$(rad) 0.54 $\mathcal{U}(0,\pi)$
$t_{c}$(s) $1126259462$ ($\equiv t_{0}$) $\mathcal{U}(t_{0}-2,t_{0}+2)$
$\phi_{c}$(rad) 4.205 $\mathcal{U}(0,2\pi)$
$e_{\text{eob}}$0.05, 0.06, 0.07, 0.08, 0.09, 0.10} $10^{-4}$
$f_{\text{ref}}$(Hz) 20 20
Table 1: Injection values of the binary parameters and corresponding priors used for PE with Bilby Ashton et al. (2019). $\mathcal{U}(a,b)$ stands for uniform prior in the range $(a,b)$. $\mathcal{\sin}(0,\pi)$ and $\mathcal{\cos}(-\pi/2,\pi/2)$ stand for the Sine and Cosine priors, respectively on the ranges mentioned. Single values represent Dirac delta priors peaked at those values. Different values in the the curly braces represent different injections. The table, therefore, represents 12 injections – 6 different eccentricities each with 2 different spin configurations.
The results of these IMRCTs are shown in Fig. 1. The top row shows the results for $(\chi_{1},\chi_{2})=(0.4,0.3)$ and the bottom row shows the same for $(\chi_{1},\chi_{2})=(0.2,0.1)$. In each of these rows, the left panel shows the iso-probability contours of the 2D probability distribution on the final mass deviation $\Delta M_{f}/\bar{M}_{f}$ (along $x$-axis) and the final spin deviation $\Delta\chi_{f}/\bar{\chi}_{f}$ (along $y$-axis) plane at 90% confidence. The right panel shows the same at 68% confidence. Different contours in each panel correspond to the injected eccentricities $e_{\text{eob}}\in$0.05, 0.06, 0.07, 0.08, 0.09, 0.1} as noted on the top of the right panel. Fig. 1 also shows the marginalized 1D posterior of $\Delta M_{f}/\bar{M}_{f}$ on top of the left panel and that of $\Delta\chi_{f}/\bar{\chi}_{f}$ on the right of the right panel.
It can be noted that in the case of moderate spin $(\chi_{1},\chi_{2})=(0.4,0.3)$, for smaller eccentricities ($e_{\text{eob}}<0.08$), the contours at 90% and 68% contain the point (0, 0) denoted by $+$ which stands for zero-deviation from GR. We will denote this point of zero-deviation from GR as the ‘GR value’. However, for larger eccentricities ($e_{\text{eob}}\gtrsim 0.08$), the contours are shifted towards positive deviation in $\Delta M_{f}/\bar{M}_{f}$ and the GR value is excluded at 68% confidence for $e_{\text{eob}}\geq 0.09$. For smaller spins $(\chi_{1},\chi_{2})=(0.2,1)$, the bias is more severe. For the low spin case (bottom row in Fig. 1), the GR value is excluded at 68% confidence for $e_{\text{eob}}=0.05$ and at 90% confidence for $e_{\text{eob}}=0.07$.
Figure 2: GR Quantile $Q_{\texttt{GR}}$ vs injected eccentricity for different component spins for the aligned spin systems. Small $Q_{\texttt{GR}}$ indicates better agreement with GR. Circles denote the quasicircular recovery and the ellipses denote the eccentric recovery. The low spin $(\chi_{1},\chi_{2})=(0.2,0.1)$ and moderate spin $(\chi_{1},\chi_{2})=(0.4,0.3)$ case is denoted by violet and orange, respectively. When recovered with quasicircular templates, low spin cases show more bias compared to moderate spin case.
One can also construct a statistical measure of the deviation from GR using the GR quantile $Q_{\texttt{GR}}$ which is the fraction of the posterior samples contained within the iso-probability contour that passes through the GR value (no-deviation point $(0,0)$). A smaller $Q_{\texttt{GR}}$ implies a better agreement with GR since the GR value is contained within a smaller iso-probability contour. In this work, we compute the GR quantile in a slightly different way. Instead of using the samples directly, we use an interpolated histogram of the 2D probability distribution. In particular, we use the implementation in summarytgr within PESummary Hoy and Raymond (2021).
In Fig. 2, we plot the $Q_{\texttt{GR}}$ for different component spin configurations as a function of the injection eccentricities. It shows the overall trend of $Q_{\texttt{GR}}$ increasing with $e_{\text{eob}}$. While for moderate spins $(\chi_{1},\chi_{2})=(0.4,0.3)$, $Q_{\texttt{GR}}$ becomes $\approx 68\%$ at $e_{\text{eob}}=0.08$, for low spins $(\chi_{1},\chi_{2})=(0.2,0.1)$, $Q_{\texttt{GR}}$ becomes $>68\%$ at $e_{\text{eob}}=0.05$ and $\gtrsim 90\%$ at $e_{\text{eob}}=0.07$.
III.1.2 Antialigned spin case
Figure 3: IMRCT with quasicircular recovery template for injections with antialigned spin configurations.The panels from the top to bottom correspond to primary antialigned, secondary antialigned and both antialigned spin configurations, respectively. For all the three cases, the dimensionless component spin magnitudes are $(\chi_{1},\chi_{2})=(0.4,0.3)$. Due to neglecting eccentricity in the recovery template, the GR value is excluded at 68% confidence at $e_{\text{eob}}=0.07,0.09$ and $0.08$ for primary antialigned, secondary antialigned and both antialigned cases, respectively. The GR value is also excluded at 90% confidence for secondary antialigned at $e_{\text{eob}}=0.1$.
In this section, we study how the bias in the IMRCT due to neglecting eccentricity in the recovery waveform model depends on the orientation of the spin with respect to the angular momentum of the system. Let the pair $(\theta^{\mathbf{S}\mathbf{L}}_{1},\theta^{\mathbf{S}\mathbf{L}}_{2})$ denote the tilts of the spin $\mathbf{S}_{1}$ of the primary and the spin $\mathbf{S}_{2}$ of the secondary, respectively, with respect to the orbital angular momentum $\mathbf{L}$ of the system. We consider the following three cases
  1. 1.
    $(\pi,0)$: $\mathbf{S}_{1}$ is antialigned and $\mathbf{S}_{2}$ is aligned.
  2. 2.
    $(0,\pi)$: $\mathbf{S}_{1}$ is aligned and $\mathbf{S}_{2}$ is antialigned.
  3. 3.
    $(\pi,\pi)$: Both $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are antialigned.
We consider only the moderate spin case where the dimensionless spins are (0.4, 0.3). The other parameters remain the same as in Table 1. We use a Dirac delta prior peaked at the injected values for the tilts. The number of injections for each of these tilts is 6 for six different eccentricities making a total of 18 injections for these three tilt configurations.
The rows from the top to bottom in Fig. 3 show the IMRCT results for the cases of primary antialigned, secondary antialigned and both antialigned, respectively. In each of these figures, the left panel shows the iso-probability contours at 90% confidence and the right panel shows the same at 68% confidence in the $\Delta M_{f}/\bar{M}_{f}-\Delta\chi_{f}/\bar{\chi}_{f}$ plane. The general trend is similar to that of aligned spin cases in Sec. III.1.1.
Dimensionless component spins Tilts Threshold eccentricity for GR violation at $90\%$ Threshold eccentricity for GR violation $68\%$
$(\chi_{1},\chi_{2})$ $(\theta^{\mathbf{S}\mathbf{L}}_{1},\theta^{\mathbf{S}\mathbf{L}}_{2})$ $e_{\text{eob}}$ at $f_{\text{ref}}=20$Hz $e_{\text{gw}}$ at $\langle f_{\text{ref}}\rangle=25$Hz $e_{\text{eob}}$ at $f_{\text{ref}}=20$Hz $e_{\text{gw}}$ at $\langle f_{\text{ref}}\rangle=25$Hz
(0.4, 0.3) (0, 0) 0.09 0.071
(0.2, 0.1) (0, 0) 0.07 0.055 0.05 0.040
(0.4, 0.3) $(\pi,0)$ 0.07 0.056
(0.4, 0.3) $(0,\pi)$ 0.10 0.079 0.09 0.071
(0.4, 0.3) $(\pi,\pi)$ 0.08 0.064
Table 2: Summary of the study of bias in IMRCT due to neglecting eccentricity. It provides the threshold injection eccentricity ($e_{\text{eob}}$ at 20 Hz) at or above which the IMRCT result is biased at 68% and 90% confidence. It also provides the eccentricity $e_{\text{gw}}$ in terms of standardized definition Shaikh et al. (2023) at an orbital averaged frequency $\langle f_{\text{ref}}\rangle=25$ Hz corresponding to each of these threshold injections. As an example, for an aligned low spin $(\chi_{1},\chi_{2})=(0.2,1.0)$ system, GR is excluded at 68% confidence for $e_{\text{eob}}=0.05$ ($e_{\text{gw}}=0.04$) and GR is excluded at 90% confidence for $e_{\text{eob}}=0.07$ ($e_{\text{gw}}=0.055$).
III.2 IMRCT with eccentric signal
In the previous section, we showed that IMRCT shows spurious deviation from GR when the parameters from an eccentric gravitational wave signal is inferred using a quasicircular waveform model (using a delta function prior on eccentricity fixed at $e_{\text{eob}}=10^{-4}$). In this section, we show that such spurious deviations go away when eccentric waveform model is used instead of quasicircular model by allowing the sampler to sample from a uniform prior on eccentricity as described in Tab. 3.
Parameter Injected value Prior
$e_{\text{eob}}$0.05, 0.06, 0.07, 0.08, 0.09, 0.10} $\mathcal{U}(8\times 10^{-3},0.15)$
$f_{\text{ref}}$(Hz) 20 20
Table 3: Injection values of the eccentricities and priors for IMRCTs with eccentric recovery.
III.2.1 Aligned spin case
Figure 4: IMRCT with eccentric recovery template for aligned spin injections. The top row corresponds to $(\chi_{1},\chi_{2})=(0.4,0.3)$ and the bottom row corresponds to $(\chi_{1},\chi_{2})=(0.2,0.1)$. Same as Fig. 1, but now the injections are recovered with eccentric templates instead of quasicircular templates. GR value is included within the iso-probability contours at both the 68% and 90% confidence intervals for all injections for these two spin configurations.
First, we consider the case of aligned spin where both the primary and the secondary spin are aligned with the angular momentum of the system, i.e., $\theta^{\mathbf{S}\mathbf{L}}_{1,2}=0$. In Fig. 4, we show the results of IMRCTs with a set of eccentric aligned spin signals. The top row shows the result for injections with moderate component spins $(\chi_{1},\chi_{2})=(0.4,0.3)$ and the bottom row shows the result for low component spins $(\chi_{1},\chi_{2})=(0.2,0.1)$. The set of injected eccentricities and the prior on the eccentricity for both of these spin configurations are provided in Tab. 3. The other parameters remain the same as in Sec. III.1.1.
In each row of Fig. 4, the left (right) panel shows the iso-probability contours of the 2D probability distribution in the $\Delta M_{f}/\bar{M}_{f}-\Delta\chi_{f}/\bar{\chi}_{f}$ plane at 90% (68%) confidence. The 1D posterior on $\Delta M_{f}/\bar{M}_{f}$ is shown on top of left panel and the 1D posterior on $\Delta\chi_{f}/\bar{\chi}_{f}$ is shown on the right of the right plot. In both of these spin configurations, for all the injected eccentricities, we find IMRCTs to be unbiased both at 68% and 90% confidence.
For these two aligned spin configurations, the GR quantile $Q_{\texttt{GR}}$, when recovered with eccentric waveform template, is shown using ellipses in Fig. 2. $Q_{\texttt{GR}}$ is plotted as a function of the injection eccentricities. It can be seen that for all the injected eccentricities, the $Q_{\texttt{GR}}$ remains $<68\%$ for low spins $(\chi_{1},\chi_{2})=(0.2,0.1)$ and $<30\%$ for moderate spins $(\chi_{1},\chi_{2})=(0.4,0.3)$ implying an agreement with GR.
III.2.2 Antialigned spin case
Figure 5: IMRCT with eccentric recovery template for injections with antialigned spin configurations. The rows from the top to bottom correspond to primary antialigned, secondary antialigned and both antialigned spin configurations, respectively. For all the three cases, the dimensionless component spin magnitudes are $(\chi_{1},\chi_{2})=(0.4,0.3)$. Same as Fig. 3 but now the injections are recovered with eccentric templates instead of quasicircular templates. Note that the GR value is included at both 68% and 90% confidence for all the injections for these three configurations.
We now turn our focus to the potential effects of the orientation of the component spins on IMRCTs with eccentric signal when recovered with eccentric waveform models. We consider the same injections as in Sec. III.1.2; we consider three possible configurations where the pair of tilts may take the values $(\pi,0),(0,\pi)$ or $(\pi,\pi)$ with the component dimensionless spins $(\chi_{1},\chi_{2})=(0.4,0.3)$. In Sec. III.1.2, these injections were recovered using a quasicircular template which resulted in spurious bias in the IMRCT result. Now we recover these injections with eccentric template using a prior on eccentricity as shown in Tab. 3.
The rows from the top to bottom in Fig. 5 shows the IMRCT contours for primary antialigned, secondary antialigned and both antialigned configurations, respectively. For a given row, the left (right) panel shows the iso-probability contours of the 2D probability distribution in the $\Delta M_{f}/\bar{M}_{f}-\Delta\chi_{f}/\bar{\chi}_{f}$ plane at 90% (68%) confidence. The 1D posterior on $\Delta M_{f}/\bar{M}_{f}$ is shown on top of the left panel and the 1D posterior on $\Delta\chi_{f}/\bar{\chi}_{f}$ is shown on the right of the right panel. Different contours in each panel correspond to different injected eccentricity as denoted on top of the right panel of each row. Unlike the results in Fig. 3, all the contours in Fig. 5 include the GR value.
We also plot the $Q_{\texttt{GR}}$ (denoted by ellipses) values of the IMRCTs corresponding to these injections in Fig. 6. Note that the $Q_{\texttt{GR}}$ values for these cases of recovery with eccentric templates are very small compared to the quasicircular recovery (denoted by circles). The $Q_{\texttt{GR}}$ for eccentric recovery is $<68\%$ for all considered eccentricities in these three antialigned spin configurations. While for the primary antialigned cases, the $Q_{\texttt{GR}}$ remains $\lesssim 10\%$, for the secondary antialigned and both antialigned cases, the GR quantile is even smaller, $Q_{\texttt{GR}}\lesssim 5\%$ implying excellent agreement with GR.
Figure 6: GR Quantile $Q_{\texttt{GR}}$ vs injected eccentricity for different component spin alignments for $(\chi_{1},\chi_{2})=(0.4,0.3)$. Small $Q_{\texttt{GR}}$ indicates better agreement with GR. Circles denote the quasicircular recovery and the ellipses denote the eccentric recovery. When recovered with quasicircular template, the $Q_{\texttt{GR}}$ tend to increase with increasing eccentricity implying increasing bias in IMRCT due to neglecting eccentricity in the waveform model. On the other hand, for eccentric recovery, the $Q_{\texttt{GR}}$ remains within $68\%$ for all eccentricities. For all cases of eccentric recovery, the $Q_{\texttt{GR}}$ is $\lesssim 10\%$. In the secondary antialigned and both antialigned cases, the $Q_{\texttt{GR}}$ becomes even smaller ($<5\%$) implying excellent agreement with GR.
IV Summary and Outlook
The IMRCT Ghosh et al. (2016, 2018) is one among several tests of GR with GWs used by the LVK. While model-agnostic in the sense that it is not based on any particular alternative theory of gravity, it does nevertheless crucially rely on the CBC template waveform being an accurate representation of the GW signal. Thus, intuitively, one would expect spurious violations of this test when physical and environmental effects are not accounted for in the waveform model.
To demonstrate this, in this work, we focus our attention on GWs from eccentric compact binaries. We inject, in zero-noise assuming an O4-like noise PSD, synthetic GW150914-like CBC signals but with a range of residual eccentricities at a reference frequency of $f_{\mathrm{ref}}=20$ Hz. The spins are always assumed to lie along the orbital angular momentum ($z-$) axis, although we consider two different sets of spin magnitudes, and all possible orientation(aligned/antialigned)-combinations of the binary components along the $z-$ axis. We then perform the IMRCT with quasicircular templates (these being routinely used by current implementations of the IMRCT), as well as eccentric templates.
We find that, as expected, the IMRCT gets progressively biased with increasing injected eccentricity while recovering with quasicircular templates. At $68\%$ confidence, we find that the IMRCT starts to get violated for eccentricities – as defined by the waveform model used – between $e_{\text{eob}}\sim 0.05-0.09$ at a reference frequency $f_{\text{ref}}=20$ Hz, for our chosen CBC systems. Converting this to a standardized definition of eccentricity Shaikh et al. (2023), these values become $e_{\text{gw}}\sim 0.04-0.07$ at an orbital averaged frequency $\langle f_{\text{ref}}\rangle=25$ Hz. Indeed, as the eccentricity is increased, the violations occur even at $90\%$ confidence.
On the other hand, when recovering with eccentric templates, we find that, the IMRCT is always satisfied, regardless of the spin magnitude, orientation, or eccentricity value, at $68\%$ confidence. Our work therefore demonstrates the need for accurate eccentric waveform models for IMRCT and possibly other tests of GR. Moreover, our work complements and expands on previous work studying the effect of neglect of eccentricity on the IMRCT. These either use approximate GW PE likelihoods Bhat et al. (2023), or only study the potential bias in the IMRCT caused by recovering a few eccentric numerical relativity waveforms with phenomenological waveforms Narayan et al. (2023).
Two important limitations of our work must be pointed out. The first is that the injections are performed in zero-noise rather than real-noise. This could potentially introduce additional fluctuations in the IMRCT as a consequence of a particular noise realisation. The second is that the recovery of the IMRCT has been tested using identical injection and recovery waveform models. It would be of interest to consider eccentric numerical-relativity injections recovered with other eccentric waveform models that are nominally used for GW PE. Such a study might potentially demonstrate non-trivial biases in the IMRCT even when eccentric waveforms are used, and if not, would ascertain the feasibility of using current eccentric waveform models for tests of GR. We plan to address the said limitations in future work.
Acknowledgements.
We thank the anonymous referees for their useful comments and suggestions. We thank Pankaj Saini for a careful reading of the draft. We also thank Prayush Kumar, Nathan K. Johnson-McDaniel, Aditya Vijaykumar, and members of the ICTS-TIFR Astrophysical Relativity Group, for useful discussions. M.A.S.’s research was supported by the Department of Atomic Energy, Government of India under project no. RTI4001 and the National Research Foundation of Korea under grant No. NRF-2021R1A2C2012473. S.A.B. acknowledges support from the Department of Science and Technology and the Science and Engineering Research Board (SERB) of India via Swarnajayanti Fellowship Grant No. DST/SJF/PSA-01/2017-18 and support from Infosys foundation. The authors acknowledge the public code at https://github.com/osheamonn/eobresum_bilby/, a modified version of which was used as the source model for injection study with Bilby Ashton et al. (2019). The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. The authors also acknowledge the use of the Alice cluster at the International Centre for Theoretical Sciences.